OpenStudy (anonymous):

find the ordered pair that represents the vector from b(-6 2) to c (-12,7) i think its bc=(-6,5)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

so -6,5?

OpenStudy (anonymous):

That's right I think.

OpenStudy (anonymous):

u sure?

OpenStudy (sleepyjess):

Let me see, this is what I am working on right now ^_^

OpenStudy (sleepyjess):

Eh, directrix is probably better at this than I am anyways :P

Directrix (directrix):

@sleepyjess I used this: <x2-x1, y2 -y1>, and got <+6,5>. No, I am not better. I am strictly cookbook on this and lacking intuition. Please proceed.

OpenStudy (sleepyjess):

I was thinking to use \(\sf < q_1 - p_1,~q_2 - p_2> = <v_1,~v_2>\) but I'm not completely sure.

Directrix (directrix):

@RadEn wrote this two years ago about this problem: vector BC = c - b = (-12-(-6), 7-2) = (... , ...)

OpenStudy (sleepyjess):

Yep, I would say your/his method is correct. I was actually just looking at that post to see if the question had been asked before. :)

Directrix (directrix):

@sleepyjess What does your version of the formula give; <6,5> or <-6,5>

OpenStudy (sleepyjess):

< -6, 5 >

Directrix (directrix):

Where is the OP?

OpenStudy (sleepyjess):

asking another question

Directrix (directrix):

I think this is correct. After all, you have studied this recently. I haven't and when I did, there was a parallelogram in the mix.

Directrix (directrix):

@cutegirl You left us.

OpenStudy (sleepyjess):

My lessons say that is the component form, so I am not sure if that is what the question wants or not?

OpenStudy (anonymous):

i here guys haha